Hi. My problem is with graphing the following equation:
f(x)=x^2
My teacher told me the f is the y of the problem. Plz help me. Thx in advance

Answer by atif.muhammad(135)

It is. f(x) is a fancy way of writing 'y'.
f(x) = x ^2 is the same as y = x^2.